OpenStudy (helpblahblahblah):

MEDAL+FAN Triangle W E T is similar to Triangle D O G. If the length of side W E is 6 feet, the length of side O G is twelve feet, and the length of side E T is 3 feet, what is the length of side D O? (hint: Make a sketch to help.) A. 1.5 feet B. 6 feet C. 15 feet D. 24 feet

OpenStudy (anonymous):



OpenStudy (anonymous):

Since \[\huge \triangle W E T \sim \triangle D O G\] hence, \[\huge \frac{WE}{DO} = \frac{ET}{OG} \rightarrow \frac{6}{DO} = \frac{3}{12}\] \[\huge \frac{6}{DO} = \frac{1}{4} \rightarrow DO = 6 \times 4 = 24 feet\] Hence option D. 24 feet is the correct answer.

OpenStudy (helpblahblahblah):

ohhhhh thanks sooo much